# On the Identifiability of Tensor Ranks via Prior Predictive Matching

### v1

@@ III A Cautionary Tale: The Tucker Model
III.1 Model Definition and Its Moments
Definition III.1 (The Tucker model).
@@ III.1 Model Definition and Its Moments
This model, introduced by Tucker (1966), is widely used for its interpretability and versatibility in multi-way analysis.
The ranks to be estimated $(r_{1},\dots,r_{M})$ are the dimensions of the core tensor $G$. As per Definition II.1 the elements of $G$ and each factor $\theta^{(p)}$ are i.i.d. random variables with moments $(\mu_{G},\sigma_{G}^{2})$ and $(\mu_{p},\sigma_{p}^{2})$, respectively.
Lemma III.2 (Moment Structure of the Tucker Model).
For the order-M Tucker model, the squared mean and the pure interaction terms involving the core ($v_{G}$), a single factor ($v_{p}$), or both ($v_{G,p}$) obey:
@@
III.2 Rank Identifiability Analysis
@@
Lemma III.3 (An Identity Among Observables).
@@
We show that both sides of the equation reduce to the same expression. Using the monomials from Lemma III.2, the left-hand side is:
@@
Theorem III.4 (Non-identifiability of Tucker ranks).
@@
We prove that the reduced design matrix $\mathbf{A}_{\text{red}}$ is the zero matrix. The identity established in Lemma III.3 provides $M$ linear dependencies among the rows of the full design matrix $\mathbf{A}$. Because these $M$ identities express every standard-basis vector $\mathbf{e}_{p}$ (the column that selects $\log r_{p}$) as a linear combination of columns that are free of any log-rank variable, the $M$ columns corresponding to the log-ranks are redundant. Consequently, the reduced design matrix $\mathbf{A}_{\text{red}}$ of Proposition II.5 is the $0\times M$ zero matrix, and the ranks are unidentifiable. ∎
IV Exemplary Tales: Identifiable Ranks with Closed-Form Expression
In contrast to the previous section, we present in the following the general solutions for the ranks of the PARAFAC/CP, Tensor Train, and Tensor Ring models for an arbitrary order $M$ tensor, showcasing how to establish identifiability according to the principles of Proposition II.5.
IV.1 PARAFAC/CP Model
@@
IV.1.1 Model Definition and Its Moments
Definition IV.1 (The PARAFAC/CP model).
@@ IV.1.1 Model Definition and Its Moments
The PARAFAC/CP decomposition, originally proposed by Harshman et al. (1970) and Carroll and Chang (1970), assumes a sum of rank-1 tensors and is widely used in multiple fields given its simple structure and applicability. The variance calculation of the CP model results in a simple and regular form for its pure interaction terms, as formalized below, a fundamental property leading to a closed-form expression for the rank $r$.
Lemma IV.2 (Monomial Moment structure of PARAFAC/CP).
@@
IV.1.2 Rank Identifiability Analysis
@@
Theorem IV.3 (Closed-Form Rank Estimator for PARAFAC/CP).
@@
We substitute the monomials from Lemma IV.2 into the expression:
@@
IV.2 The Tensor Train Model
@@
IV.2.1 Model Definition and Its Moments
Definition IV.4 (The Tensor-Train model).
@@ IV.2.1 Model Definition and Its Moments
The ranks to be estimated are the internal bond dimensions $(r_{1},\dots,r_{M-1})$, which we refer to as interior ranks. A full derivation of the general TT moments involves a series of (repeted) algebraic calculations, which we present in Appendix A.3.
@@
Lemma IV.5 (Monomial Moment Structure of Tensor-Train).
@@
IV.2.2 Rank Identifiability Analysis
@@
Theorem IV.6 (Closed-Form Interior-Rank Estimator for TT).
@@
The proof follows by dividing the two identities from Lemma IV.5. The unknown coefficient of variation term, $\sigma_{p}^{2}/\mu_{p}^{2}$, cancels, resulting in the rank equation:
@@
IV.3 The Tensor Ring (TR) Model
The Tensor Ring (TR) decomposition generalizes the Tensor Train by imposing cyclic boundary conditions that connect the first and last latent dimensions, closing the chain into a ring. Formally, for an order-$M$ tensor with TR ranks $(r_{1},\dots,r_{M})$ and indices modulo $M$, the mean rate tensor is defined as
where each core matrix $\Theta^{(p)}_{i_{p}}\in\mathbb{R}^{r_{p-1}\times r_{p}}$ and the boundary condition $r_{0}=r_{M}$ closes the loop. All entries of $\Theta^{(p)}_{i_{p}}$ are i.i.d. with moments $(\mu_{p},\sigma_{p}^{2})$.
IV.3.1 Moment structure
The TR model preserves the factorized moment structure of the TT and Tucker models while introducing new constraints from its cyclic topology. In particular, the first and second prior predictive moments yield the monomials
$(E[Y])^{2}$ $=\Big(\prod_{k=1}^{M}r_{k}\mu_{k}\Big)^{\!2}$ (7)
$v_{\{p,p+1\}}$ $=\Big(\prod_{k=1}^{M}r_{k}\Big)\sigma_{p}^{2}\sigma_{p+1}^{2}\!\!\!\!\!\prod_{k\notin\{p,p+1\}}\!\!\!\!\!\mu_{k}^{2}$ (8)
whose asymmetric exponents ($r_{p-1}^{1}$ vs. $r_{p}^{2}$) reflect the directed nature of the incoming and outgoing links at each variance node. A full derivation with explicit double-counting verification appears in Appendix A.4.2.
IV.3.2 Rank Identifiability Analysis
Combining these three observable moment families produces a simple, exactly-cancelling statistic:
$\xi_{p}:=\frac{(E[Y])^{2}\,v_{\{p,p+1\}}}{v_{\{p\}}\,v_{\{p+1\}}},\qquad\Rightarrow\qquad\xi_{p}=r_{p}\,\frac{r_{p-1}}{r_{p+1}}.$ (10)
This relation—proved in Appendix A.4.2—encodes all rank dependencies implied by the cyclic topology. Taking logarithms, $x_{p}=\log r_{p}$, $\psi_{p}=\log\xi_{p}$, yields the linear circulant system
$x_{p}+x_{p-1}-x_{p+1}=\psi_{p},\qquad p=1,\dots,M.$
Closed-form recovery and identifiability.
The coefficient matrix of the above system is circulant with first row $(1,-1,0,\dots,0,1)$. Its discrete Fourier eigenvalues are $\lambda_{k}=1+2i\sin(2\pi k/M)$, all nonzero for $M\geq 3$, implying full-rank invertibility and hence identifiability of the TR ranks. The solution has a simple closed form in the Fourier domain:
$\widehat{x}_{k}=\widehat{\psi}_{k}/\lambda_{k},\qquad x=\mathrm{DFT}^{-1}(\widehat{x}),\qquad r_{p}=\exp(x_{p}).$
Edge cases follow the expected pattern: $M=2$ reduces to a matrix/CP model, and $M=3$ remains identifiable (eigenvalues $1,\,1\pm i\sqrt{3}$).
The TR model thus shares the clean identifiability property of the TT model, despite its cyclic symmetry. The key difference is that the ring topology leads to a first-order circulant system whose unique invertibility ensures all ranks $(r_{1},\ldots,r_{M})$ are determined directly from first and second moments. All intermediate algebraic steps and the detailed counting argument are given in Appendix A.4.2.
V The Tensor-Rank Estimation Pipeline
We now present a complete, algorithmically specified, robust pipeline for applying the theoretical results of this work. It consists of estimating from tensor-data the pure-interaction terms $\{\hat{v}_{S}\}$ and the estimated mean $\hat{E}[Y]$, to replace them in the specific closed-form expression derived in Section IV for the ranks of each identifiable model. Recall that all rank estimation expressions are written as a division, i.e., the rank $r_{p}$ is always calculated via an equation of the $\frac{\text{Num}_{p}}{\text{Den}_{p}}$ form, with $\text{Num}_{p}$ the numerator and $\text{Den}_{p}$ the denominator, with expressions adjusted for each specific model.
@@ V The Tensor-Rank Estimation Pipeline
    Total Covariances: For every non-empty pattern of shared indices $S\subseteq\{1,\dots,M\}$, compute the sample covariance $\hat{C}_{S}$ by drawing a (large) number of pairs of entries $(\mathbf{a}_{j},\mathbf{b}_{j})$ that match exactly on the index set $S$:
@@
To ensure robustness and quantify the estimation uncertainty, we repeat the moment estimation procedure above on bootstrap replicates of the observed tensor-data. For $b=1,\dots,B$:
@@
    These bootstrap moments are plugged into the appropriate closed-form formula described in Section IV to yield the numerator $\text{Num}_{p}^{(b)}$ and denominator $\text{Den}_{p}^{(b)}$ for each bootstrapped replicate of the identifiable rank $r_{p}^{(b)}$:
@@
    Additionally, each bootstrap estimate is regularized to prevent numerical instability. For each replicate $b$, we propose a regularized rank
@@
We provide point rank estimates with bootstraped confidence intervals based on the collection of regularized estimates $\{\hat{r}_{p}^{(b)}\}_{b=1}^{B}$:
@@
VI Empirical Validation
We validate the theoretical findings for the identifiable models using simulation studies, confirming that the closed-form estimators accurately recover the true latent ranks when applied on finite data.
VI.1 Simulation Setup
@@
We generated synthetic tensor-data according to the PARAFAC/CP. The elements of the latent factors ($\theta^{(p)}$) for both models were drawn independently from Gamma distributions. The shape and scale parameters of the Gamma priors were chosen to control the factor-specific means and variances ($\mu_{p},\sigma_{p}^{2}$) while ensuring the resulting rate tensor $\bm{\eta}$ was strictly positive. The final observed tensor elements $Y_{\mathbf{i}}$ were then drawn from a Poisson distribution with the corresponding rate $\eta_{\mathbf{i}}$:
$Y_{\mathbf{i}}\sim\text{Poisson}(\eta_{\mathbf{i}})\;.$
@@
We study the precision of the constructed pipeline by generating data in a wide range of conditions:
  • For PARAFAC/CP, varying the tensor order ($M\in\{3,4\}$), the tensor dimensions ($N_{p}\in\{(100,100,100),(50,50,50,50)\}$), and the ground-truth latent ranks ($r_{p}\in\{5,\ldots,65\}$).
For each case, we aggregate results over multiple independent runs, i.e., realizations of both the data and the latents using a set of values for the prior described in Appendix C. For each run, a median rank was estimated from a distribution of bootstrap replicates, and the plots show the mean of these median estimates.
VI.2 Results for Identifiable Models
PARAFAC/CP Model.
The results for the PARAFAC/CP model provide strong empirical support for its theoretical identifiability. Figure 1 shows the estimated CP rank versus the true CP rank, aggregated across all simulation runs.
@@ VI.2 Results for Identifiable Models
The mean of the estimated ranks (green dots) closely tracks the line of perfect estimation. For true ranks ranging from low values up to 90, the estimator correctly identifies the latent dimension and follows the linear trend. Best performance (both in closer mean and reduced estimation uncertainty) is achieved for reduced (<25<25) rank values.
@@
We observe a slight positive bias and an increase in estimation variance (error bars) as the true rank grows. This is expected, as higher-rank structures are more complex, and their corresponding moment-based signatures are more sensitive to the statistical noise inherent in finite data sampling.
Overall, the results confirm that the estimator is effective and that the CP rank is practically identifiable.
Further empirical studies for TT and TR are presented in the Appendix C.
VII Discussion
@@
Our prior predictive moment matching framework reveals that rank identifiability in probabilistic tensor models hinges on the latent factors’ interaction topology. More precisely, we demonstrate that linear (additive) structures like PARAFAC/CP yield monomial moments with clean algebraic cancellations, enabling closed-form estimators. Chain-like topologies (TT) introduce asymmetry via contiguous/non-contiguous interactions, providing independent equations per rank that also enable their empirical estimation. The cyclic structure in TR maintains solvability through invertible circulant matrices, despite its symmetry. In contrast, multiplicative “hub-and-spoke” topologies (e.g., Tucker) induce linear dependencies, rendering the reduced system degenerate – we especulate that higher moments or correlations structure in the core tensor could render the models’ ranks identifiable.
@@
This topology-based lens offers a principled tensor model analysis toolbox: for identifiable models, low-order moments suffice for exact rank recovery; for others, heuristics like matricization (see Appendix A.2) implicitly reduce to CP-like problems.
Limitations of the presented work include the i.i.d. prior assumption, and the focus on second-order moments only —we hypothesize extensions to higher-order moments or correlated priors could broaden its applicability. Empirically, our pipeline is robust to moderate noise, but sensitive to extreme sparsity which requires further numerical considerations.
@@
A natural next-step is the application of the proposed framework to real-data sets (e.g., on fMRI tensors). More ambitious future work might consider integration of the prior predictive moment matching approach into variational inference for scalable tensor ML, or exploring non-Euclidean topologies (e.g., hyperbolic embeddings). A natural generalization of our framework is to consider generic tensor factorizations defined by arbitrary tensor networks (Cichocki et al., 2016). In this case, the latent ranks would correspond to the bond dimensions along network edges, and identifiability would depend on the network’s topology (e.g., tree-like vs. loopy structures).

### v2

@@ 3 A CAUTIONARY TALE: THE TUCKER MODEL
3.1 Model Definition and Its Moments
Definition 3.1 (The Tucker model).
@@ 3.1 Model Definition and Its Moments
This model, introduced by Tucker (1966), is widely used for its interpretability and versatility in multi-way analysis.
The ranks to be estimated $(r_{1},\dots,r_{M})$ are the dimensions of the core tensor $G$. As per Definition 2.1 the elements of $G$ and each factor $\theta^{(p)}$ are i.i.d. random variables with moments $(\mu_{G},\sigma_{G}^{2})$ and $(\mu_{p},\sigma_{p}^{2})$, respectively.
Lemma 3.2 (Moment Structure of the Tucker Model).
For the order-M Tucker model, the squared mean and the pure interaction terms involving the core ($v_{G}$), a single factor ($v_{p}$), or both ($v_{G,p}$) obey —see detailed derivations in Appendix A.2:
@@
3.2 Rank Identifiability Analysis
@@
Lemma 3.3 (An Identity Among Observables).
@@
We show that both sides of the equation reduce to the same expression. Using the monomials from Lemma 3.2, the left-hand side is:
@@
Theorem 3.4 (Non-identifiability of Tucker ranks).
@@
We prove that the reduced design matrix $\mathbf{A}_{\text{red}}$ is the zero matrix.
The identity established in Lemma 3.3 provides $M$ linear dependencies among the rows of the full design matrix $\mathbf{A}$. Because these $M$ identities express every standard-basis vector $\mathbf{e}_{p}$ (the column that selects $\log r_{p}$) as a linear combination of columns that are free of any log-rank variable, the $M$ columns corresponding to the log-ranks are redundant. Consequently, the reduced design matrix $\mathbf{A}_{\text{red}}$ of Proposition 2.5 is the $0\times M$ zero matrix, and the ranks are unidentifiable. ∎
4 EXEMPLARY TALES: IDENTIFIABLE RANKS WITH CLOSED-FORM EXPRESSION
In contrast to the analysis in the previous section, we present below the general solutions for the ranks of the PARAFAC/CP, Tensor Train, and Tensor Ring models for an arbitrary order $M$ tensor, showcasing how to establish identifiability according to the principles formalized in Proposition 2.5.
4.1 PARAFAC/CP Model
@@
4.1.1 Model Definition and Its Moments
Definition 4.1 (The PARAFAC/CP model).
@@ 4.1.1 Model Definition and Its Moments
The PARAFAC/CP decomposition, originally proposed by Hitchcock (1927) , with later formulations by Harshman and others (1970) and Carroll and Chang (1970), assumes a sum of rank-1 tensors and is widely used in multiple fields given its simple structure and applicability. The variance calculation of the CP model results in a simple and regular form for its pure interaction terms, as formalized below, a fundamental property leading to a closed-form expression for the rank $r$.
Lemma 4.2 (Monomial Moment structure of PARAFAC/CP).
@@
4.1.2 Rank Identifiability Analysis
@@
Theorem 4.3 (Closed-Form Rank Estimator for PARAFAC/CP).
@@
We substitute the monomials from Lemma 4.2 into the expression:
@@
4.2 The Tensor Train Model
@@
4.2.1 Model Definition and Its Moments
Definition 4.4 (The Tensor-Train model).
@@ 4.2.1 Model Definition and Its Moments
The ranks to be estimated are the internal bond dimensions $(r_{1},\dots,r_{M-1})$, which we refer to as interior ranks. A full derivation of the general TT moments involves a series of (repeated) algebraic calculations, which we present in Appendix A.3.
@@
Lemma 4.5 (Monomial Moment Structure of Tensor-Train).
@@
4.2.2 Rank Identifiability Analysis
@@
Theorem 4.6 (Closed-Form Interior-Rank Estimator for TT).
@@
The proof follows by dividing the two identities from Lemma 4.5. The unknown coefficient of variation term, $\sigma_{p}^{2}/\mu_{p}^{2}$, cancels, resulting in the rank equation:
@@
Remark 4.7 (The Tensor Ring Model).
A generalization of the TT, known as the Tensor Ring (TR) model, can be formed by starting from the TT and connecting its beginning and end, creating a closed loop with no limitations on the outer rank dimension. We prove the identifiability of such TR model, and present its corresponding rank estimation formulas in Appendix  A.4.
5 THE TENSOR-RANK ESTIMATION PIPELINE
We now present a complete, algorithmically specified, robust pipeline for applying the theoretical results of this work: it consists of estimating from observed tensor-data the pure-interaction terms $\{\hat{v}_{S}\}$ and the estimated mean $\hat{E}[Y]$, to replace them in the specific closed-form expression derived in Section 4 for the ranks of each identifiable model.
Recall that all rank estimation expressions are written as a division, i.e., the rank $r_{p}$ is always calculated via an equation of the $\frac{\text{Num}_{p}}{\text{Den}_{p}}$ form, with $\text{Num}_{p}$ the numerator and $\text{Den}_{p}$ the denominator; where expressions are adjusted for each specific model.
@@ 5 THE TENSOR-RANK ESTIMATION PIPELINE
    Total Covariances: For every non-empty pattern of shared indices $S\subseteq\{1,\dots,M\}$, compute the sample covariance $\hat{C}_{S}$ by drawing a (large) number of pairs of entries $(\mathbf{a}_{j},\mathbf{b}_{j})$ that match exactly on the index set $S$, i.e.,
@@
To ensure robustness and quantify the estimation uncertainty, we repeat the inner moment estimation procedure described above on bootstrap replicates of the observed tensor-data.
For $b=1,\dots,B$:
@@
    These bootstrap moments are plugged into the appropriate closed-form formula described in Section 4 to yield the numerator $\text{Num}_{p}^{(b)}$ and denominator $\text{Den}_{p}^{(b)}$ for each bootstrapped replicate of the identifiable rank $r_{p}^{(b)}$:
@@
    Each bootstrap estimate is regularized to prevent numerical instability. For each replicate $b$, we propose a regularized rank
@@
We provide point rank estimates with bootstrapped confidence intervals based on the collection of regularized estimates $\{\hat{r}_{p}^{(b)}\}_{b=1}^{B}$:
@@
6 Empirical Validation
We validate the theoretical findings for the identifiable models using simulation studies, confirming that the closed-form estimators accurately recover the true latent ranks when applied to finite tensor data.
We present results for the PARAFAC/CP model in Section 6.1, and the empirical studies for TT and TR in Appendix C. In addition, we present in Section 6.2 a comparison of the proposed rank estimation procedure and Bayesian inference alternatives on the real-world PINCAT MRI dataset (Candes et al., 2013) [^2] Link to the dataset https://candes.su.domains/software/SURE/data.html.
The codebase with the tensor rank estimation pipeline to replicate the experiments is available in https://github.com/iurteagalab/tensor-ranks.
6.1 Simulations of the PARAFAC/CP model
@@
We generated synthetic tensor-data according to the PARAFAC/CP model, where elements of the latent factors ($\theta^{(p)}$) were drawn independently from Gamma distributions. The shape and scale parameters of the Gamma priors were chosen to control the factor-specific means and variances ($\mu_{p},\sigma_{p}^{2}$) while ensuring the resulting rate tensor $\bm{\eta}$ was strictly positive. The final observed tensor elements $Y_{\mathbf{i}}$ were then drawn from a Poisson distribution with the corresponding rate $\eta_{\mathbf{i}}$ and $Y_{\mathbf{i}}\sim\text{Poisson}(\eta_{\mathbf{i}})\;.$
@@
We study the precision of the constructed pipeline by generating PARAFAC/CP model data in a wide range of conditions: we vary the tensor order ($M\in\{3,4\}$), the tensor dimensions ($N_{p}\in\{(100,100,100),(50,50,50,50)\}$), and the ground-truth latent ranks ($r_{p}\in\{5,\ldots,65\}$).
For each case, we aggregate results over multiple independent runs, i.e., realizations of both the data and the latents, using a set of values for the prior described in Appendix C. For each run, a median rank was estimated from a distribution of bootstrap replicates, and the plots show the mean of these median estimates.
@@
We provide strong empirical support of the theoretical identifiability and estimation accuracy of the PARAFAC/CP model’s ranks. Figure 1 shows the estimated CP rank versus the true CP rank, aggregated across all simulation runs. The mean of the estimated ranks (green dots) closely tracks the line of perfect estimation. For true ranks ranging from low values up to 90, the estimator correctly identifies the latent dimension and follows the linear trend.
@@
Best performance (both in closer mean and reduced estimation uncertainty) is achieved for reduced (<25<25) rank values. We observe a slight positive bias and an increase in estimation variance (error bars) as the true rank grows. This is expected, as higher-rank structures are more complex, and their corresponding moment-based signatures are more sensitive to the statistical noise inherent in finite data sampling. Overall, the results confirm that the estimator is effective and that the CP rank is practically identifiable.
Comparison with AIC/BIC Rank Estimation.
To benchmark the efficacy of the proposed prior-predictive (PP) estimator, we evaluate it alongside information-criteria approaches (AIC/BIC) to rank estimation. Notably, the variational CP model coupled with AIC/BIC scoring rules consistently collapsed to the lowest available rank (often $R=2$) in the grid search of our experiments, across the simulated CP datasets under various prior assumptions (Gamma, Uniform, Gaussian).
In contrast, the PP estimator consistently recovered the scale of the ground-truth rank across non-zero mean scenarios. [^3] If all factors have zero-mean priors, all pure interaction terms vanish and hence, tensor ranks are not identifiable from second-order moments.
6.2 Real-world tensor-data experiment.
We additionally assess rank estimation performance on the real-world PINCAT MRI dataset, by comparing our automatically determined rank ($\hat{r}_{\text{PP}}=70$) with the rank values obtained by explicit Bayesian inference methods described by Zhang et al. (2022). Specifically, we compare our approach to the ranks obtained via BFCP and Hamiltonian Monte Carlo (HMC) methods using a hold-out of 10% of the data to measure the predictive performance.
Table 1 illustrates that our computationally inexpensive PP estimator matches the performance and uncertainty calibration of full Bayesian procedures.
Table 1: PINCAT MRI rank estimation and predictive performance against Bayesian baselines.
Method Rank Test RMSE Test SD
Prior Predictive (ours) 70 0.6086 0.4876
BFCP (Zhang et al., 2022) 17 0.6087 0.4877
HMC (Zhang et al., 2022) 65 0.6086 0.4876
Robustness to Missing Data.
A crucial aspect of real-world tensor analysis is the presence of missing entries. We stress-tested our PP estimator under Missing Completely At Random (MCAR) conditions, applying random masks incrementally (up to 50% data missingness) on the PINCAT dataset.
As the percentage of observed data decreases, the variance of the estimator increases (naturally matching the reduction in valid paired observations, $N_{\text{cov}}$). We observe in the results presented in Table 2 that the estimated rank shrinks gradually due to the reduced effective information: the PP estimator degrades smoothly and predictably, tracing the theoretical finite-sample boundaries.
Table 2: PINCAT MRI predictive and rank estimation performance under varying MCAR missingness.
Observed % Test RMSE Estimated Rank ($\hat{r}_{\text{PP}}$)
90% 0.6086 106.9
80% 0.6414 83.0
70% 0.6781 66.4
50% 0.7609 60.9
7 DISCUSSION
We present a prior predictive moment matching analysis framework revealing that rank identifiability in probabilistic tensor models hinges on the latent factors’ interaction topology.
@@ 7 DISCUSSION
More precisely, we demonstrate that linear (additive) structures like the PARAFAC/CP model yield monomial moments with clean algebraic cancellations, enabling closed-form estimators. Chain-like topologies (TT) introduce asymmetry via contiguous/non-contiguous interactions, providing independent equations per rank that also enable their empirical estimation. The cyclic structure in TR maintains solvability through invertible circulant matrices, despite its symmetry.
In contrast, multiplicative “hub-and-spoke” topologies (e.g., Tucker) induce linear dependencies, rendering the reduced system degenerate —we speculate that higher moments or correlation structure in the core tensor could render the models’ ranks identifiable.
@@
The presented topology-based lens offers a principled tensor model analysis toolbox: for identifiable models, low-order moments suffice for exact rank recovery; for others, heuristics like matricization (see Appendix A.2) implicitly reduce to CP-like problems.
Limitations of the presented work include the i.i.d. prior assumption —correlated prior-based moments remain log-linearizable, but they require an augmented parameterization, analogous to matrix results by (Zhu and Sato, 2025)— and the focus on second-order moments only —we hypothesize extensions to higher-order moments are possible and would broaden the applicability of the presented analysis.
Empirically, our pipeline is shown to be robust to moderate noise, but sensitive to extreme sparsity and data-missingness, which requires both further numerical considerations and robustness mechanisms. A natural next-step is the application of the proposed framework to other real-data sets (e.g., on fMRI tensors).
More ambitious future work might consider integration of the prior predictive moment matching approach into variational inference for scalable tensor ML, or exploring non-Euclidean topologies (e.g., hyperbolic embeddings).
@@
A natural generalization of our framework is to consider generic tensor factorizations defined by arbitrary tensor networks (Cichocki et al., 2016). In this case, the latent ranks would correspond to the bond dimensions along network edges, and identifiability would depend on the network’s topology (e.g., tree-like vs. loopy structures).